# Shape-specific fluctuations of an active colloidal interface

### v2

@@ II Model
$\dot{\mathbf{r}}_{i}$ $=v_{s}\,\mathbf{e}_{i}+\mu\,\mathbf{F}_{i}^{b}+\sqrt{2D_{t}}\,\bm{\xi}^{t}_{i},$ (1)
$\dot{\theta}_{i}$ $=\chi_{r}\left(\mathbf{e}_{i}\times\mathbf{J}_{i}\right)+\sqrt{2D_{r}}\,\bm{\xi}^{r}_{i}.$ (2)
In the above equations, $\mu$ is mobility, $D_{t}$ and $D_{r}$, respectively, are translational and rotational diffusion constants of the particle, while $\bm{\xi}^{t}$ and $\bm{\xi}^{r}$ are white noises with zero mean and unit variances. The constant $\chi_{r}$ is taken to be positive so that the particles rotate away from each other (chemo-repulsive). The force on the $i$th particle is given as: $\mathbf{F}_{i}^{b}=-{\partial U}/{\partial\mathbf{r}_{i}}$, while we have: $U=\sum^{N-1}_{i=1}\mathcal{U}^{b}(\mathbf{r}_{i},\mathbf{r}_{i+1})$. Here, $\mathcal{U}^{b}(\mathbf{r}_{i},\mathbf{r}_{j})=k\left(r_{ij}-r_{0}\right)^{2}$ is a harmonic potential of stiffness $k$ and natural length $r_{0}$ which holds the chain together, while $r_{ij}=|\mathbf{r}_{i}-\mathbf{r}_{j}|$.
@@
$\theta_{i}(0)=0,\qquad x_{i}(0)=0,\qquad\forall i,\qquad\qquad$ (4)
$y_{i+1}(0)=y_{i}(0)+2b,\qquad\,1\leq i\leq N-1$ (5)
where $b$ is the radius of each colloid. Note that $r_{0}=2b$. All numerical results in this paper are generated with an explicit Euler-Maruyama integrator with time step $dt=0.01$ and simulation time $T=7.2\times 10^{5}$. For all simulations here (unless stated otherwise) $b=1$, $c_{0}=1$, $D_{c}=1$, $\hat{\mu}=\mu k_{b}=91.6s^{-1}$, and $v_{s}=1$. $D_{r}=0$ for all sections, with the exception of Sec. V.3 where values are explicitly stated. Wherever realization average is performed, we use $50$ realizations, with the exception of $10$ used in Appendix .1.
@@ III Dynamical regimes
The dimensionless number $\Lambda$ quantifies the competition between deterministic rotations (due to phoretic interactions) and deterministic propulsion. The dimensionless number $\tilde{\Lambda}$ quantifies deterministic and random motion in the positional sector. A variant on the above model, section II, was introduced in [25] and studied further in [29]. In both these papers, we had considered the role of trail created by the particles. Here, we first assume instantaneous chemical interactions, and discuss the effect of trail-mediated history later (in Appendix .1). Additionally, we note that the previous works ignored the role of translational noise $D_{t}$. In [29], it was found that a competition between $\tau_{r}$ and the chemical diffusivity of trails leads to the formation of a stable C-shape chain.
@@
It is informative to study the model described above in terms of the two dimensionless numbers. In Fig. 1, we present a phase diagram in the plane of $\Lambda$ and $\tilde{\Lambda}$. We find that the C-shape is sustained for a selected range of dimensionless numbers $\Lambda$ and $\tilde{\Lambda}$ [25, 29]. Let us construct in addition a fluctuating length scale $l_{f}=\frac{D_{t}}{v_{s}}$ and a correlation length scale $l_{C}=\frac{\chi_{r}}{bv_{s}}$ to aid with the analysis. First, in the limit of $\Lambda\gtrsim 10^{2},\tilde{\Lambda}\gtrsim 10^{1}$, there exists insufficient rotations and the chain remains flat within the simulation time scale [29]. This instead corresponds to the region of extremely short-wavelength fluctuations ($l_{f}<b$) and small deterministic correlation lengths ($l_{C}<10^{-3}b$). This straight-chain phase is labelled as Phase I (see Fig. 1(b) top left). Next, for regions $\tilde{\Lambda}\lesssim 10^{0}$, we find a disordered phase with effectively zero positional order in the chain. Here we find that $l_{f}>b$, thus we conclude that long-wavelength fluctuations of greater than the monomer size are not supported by our colloidal chain. This disordered phase is labelled as Phase II (Fig. 1(b) top right). Further, focusing in the regime $\tilde{\Lambda}>10^{1}$ and $\Lambda<10^{-3}$ corresponds to the case where deterministic rotations are instantaneous across the system ($\tau_{f}<<\tau$), such that the chemical interactions are instantaneously correlated. Here one has that $l_{C}\gtrsim 10^{3}b$, thus correlations much larger than the system size. The dynamics of the chain displays “frustrated” behavior, not attaining any non-trivial spatial structure as there is not enough time to respond to instantaneous chemical gradients. This frustrated phase is labelled as Phase III (Fig. 1(c) bottom left). The remaining region corresponds to that of the C-shape (Phase IV). We thus conclude that the C-shape thus corresponds to regimes where the correlation lengths $l_{C}$ are of the order of the chain length, which is in addition stable to short-wavelength fluctuations (of typical distance less that one monomer radius). It is this dynamical steady-state (DSS) phase that we study in the rest of this paper.
@@
We also note that, within the C-shape phase, its existence as a DSS depends on the choice of initial conditions of Eq. (4). Deviating from these does not render a universality of the DSS, unless one includes the existence of chemical trails in the model [25, 29]. This scenario is further discussed in Appendix .1. If one deviates from the purely chemo-repulsive scenario considered here, the presence of trails are also necessary for other non-trivial DSS, such as the swimming/undulating state. However, various structural SS (e.g. designer crystallites that break handedness symmetry) is possible via the instantaneous chemical deposition model considered here, though one would have to significantly alter from the fully chemo-repulsive scenario, and incorporate mixed (anti-symmetric and/or non-reciprocal) interactions [29].
@@ III.1 Characteristics of the C-shape
    Finite curvature along chain (or the averaged local curvature)
@@
As explained in Appendix .7, there are various other DSS that meet some of the above requirements but not all. For instance the aforementioned Phase I DSS satisfies only (i), (iii) and (iv). The C-shape is unique in that it is the only DSS that meets requirements (i)-(iv), with its propulsion perpendicular to the direction of the body axis. It is in this specific phase in which the fluctuations are computed.
@@
Table 1: Table of exponents defined in equations (8-14) of section IV. Note that $z$ is defined such that $\beta^{(2)}$ is in the denominator for the first and third rows, whereas $\beta^{(1)}$ is in the denominator for the others (orientational and locally flat exponents). For the last two columns, $z$ is taken to be $z^{\prime}$ implicitly. For comparison, we note that for a one-dimensional interface: $\alpha^{\mathrm{EW}}=1/2$, $\beta^{\mathrm{EW}}=1/4$; $\alpha^{\mathrm{KPZ}}=1/2$, $\beta^{\mathrm{KPZ}}=1/3$ [34, 35].
@@ III.2 C-shape selection
In Phase IV, as mentioned previously, the C-shape is always attained under the initial conditions of 4. We note that in such an autophoretic chain with chemical trails this DSS is the universal solution regardless of initial conditions [25, 29] (see Appendix .1)). However, it is instructive to ask to what extent the C-shape is attained under different initial conditions.
To this end, we apply a hysteretic protocol by varying $\Lambda$ and carrying on the final state of a given simulation to the initial conditions of the simulation at the next time step. The order parameter used to identify this DSS is the absolute value of the curvature, $|\kappa|$. The results of these are presented in Fig. 1(c). Two separate cases are presented, corresponding to decrease of $\Lambda$ starting from Phase I (the stiff chain; Fig. 1(c), top), and increase of $\Lambda$ starting from Phase IV (the C-shape; Fig. 1(c), bottom). We observe from the former that the stiff-to-C transition is irreversible. Thus, a C-shape DSS cannot lose its spatial curvature via a decrease in self-phoretic torques. From the latter we see that the C-to-stiff transition is hence not permitted; both forward and backward sweeps of the protocol do not change the overall curvature. We also note that we have separately varied $\tilde{\Lambda}$ to and from Phase IV (not shown here); the results of these indicate that selected random configurations of Phase II may indeed transition to IV, but again there is no clear reliable reversibility of transition to define a hysteresis loop. This irreversibility of the first transition implies that such a hysteresis curve and hence mapping to equilibrium classes of phase transitions is not well defined (indeed one may have to devise an alternate order parameter); we leave the exact characterization of the nature of the transition to and from the C-shape for future work.
@@ IV Interface fluctuations
For the C-shape chain propelling in the positive x-direction, one can define a height function $h(y)$ along the chain. An example of its temporal evolution is shown in Fig. 1(d). One can thus choose a particular length segment along the chain and compute the height fluctuations, and measure its root-mean-squared
@@
The quantities of interest are both the dynamical ($W_{h}(t)$ and $W_{\theta}$) and steady-state ($W^{\mathrm{ss}}_{h}(t)$ and $W^{\mathrm{ss}}_{\theta}(t)$) properties of these fluctuations. These are:
$W_{h}(t)$ $\sim t^{\beta_{h}},\qquad W^{\mathrm{ss}}_{h}\sim N^{\alpha_{h}},$ (10)
$W_{\theta}(t)$ $\sim t^{\beta_{\theta}},\qquad W^{\mathrm{ss}}_{\theta}\sim N^{\alpha_{\theta}}.$ (11)
@@
To compute the quantities $W_{h}(t)$ and $W_{\theta}(t)$ [Eq. (8) and (9)] respectively, we discuss two possibilities. There are two separate segments about which fluctuations are measurable. Let us label the length of this segment to be $N^{\prime}$. The first obvious segment corresponds to $N^{\prime}=N$, corresponding to the average over the entire C-shape. The second corresponds to that of a locally flat region; here $N^{\prime}<<N$, which is taken to be sufficiently far from the edges of the chain. Though established results exist for fluctuations about curved segments of an interface [9], these typically correspond to that of a (subset of a) circular interface, thus different from that of the C-shape (see Appendix .5). For the latter, it is to be noted that fluctuations of growing (circular) interfaces are typically measured in this regime [36, 37] ; the scaling properties in this region are almost always equal to that of globally measured quantities - for instance in [16] for a flat surface and [36] for a curved surface. We discuss the specific instances of equivalences and/or differences for the C-shape below.
All the results presented in this paper are computed for $N$ values of $N\in[32,64,100,128,192,200,256,312]$, with curves for specific $N$ displayed (and labelled) in each respective figure. Results for the incorporation of rotational noise are computed for $N\in[64,128,192,256,312]$. Choices for $N^{\prime}$ are stated explicitly in the respective section.
@@ V Results
A summary of the results in this paper is presented in Table 1, wherein the respective Figures are referenced. It is to be noted that the value of $\beta$ for each respective figure is obtained by obtaining the best collapse of the scaling relations 12 and 14 (by eye). The values for $\alpha$ on the other hand are obtained by a least-squares fit on the steady state values of $W$ for each system size. For each plot of $W$, the fitting windows used for the in-figure $\beta$ line of best fit is mentioned in the caption.
@@ V.1 Height fluctuations
From (18) and Fig. 2, we find that regimes (iii) and (iv) exhibit a clear FV scaling with a roughness exponent $\alpha_{h}$ and a dynamic exponent $z_{h}=\frac{\alpha_{h}}{\beta_{h}}\approx 0.52$. To the best of our knowledge, this set of exponents does not correspond to any previously reported universality class. We thus interpret this as a novel “C-shape roughness” scaling, distinct to those found for either circular or flat interfaces [36, 15]. The super-ballistic exponent can be appreciated by noting that as the C-shape formation takes place, displacements of each monomer from the mean is enhanced in either direction (along the x-axis) along the chain. These squared deviations from the mean height are in addition symmetric along the y-axis. $l_{c}$ is clearly enhanced due to this symmetry. It is to be noted that in a conventional ballistic scaling one simply has uni-directional deviations along the propulsion axis; in our system this is applicable to the mean $h(y)$; clearly $l_{c}$ in this case is smaller than for a symmetric C-shape. We suggest that these combined effects give rise to the super-ballistic growth exponent $\beta_{h}$. The presence of a super-ballistic exponent can be rationalized via a linearised analysis of the EOM 2. This is presented in Appendix .3. It is to be noted that varying $\tau_{r}$ changes the time at which super-ballistic growth sets in (exponent and roughness is unchanged). The roughness exponent $\alpha_{h}$, on the other hand, indicates the relative growth of height fluctuations when $l_{c}$ spans the system size. Here, the system is is at $\sim 10^{2}\tau_{f}$ and the C-shape is fully formed. We note that the value we report is in between the standard roughness encountered in dynamics without conservation [34, 35] and smaller than models with mass conservation [38, 17, 39]. In addition larger than most numerical deposition models such as for instance the solid-on-solid growth model [16]. We discuss these discrepancies further in the Discussion section.

### v3

@@ II Model
$\dot{\mathbf{r}}_{i}$ $=v_{s}\,\mathbf{e}_{i}+\mu\,\mathbf{F}_{i}^{b}+\sqrt{2D_{t}}\,\bm{\xi}^{t}_{i},$ (1a)
$\dot{\theta}_{i}$ $=\chi_{r}\left(\mathbf{e}_{i}\times\mathbf{J}_{i}\right)+\sqrt{2D_{r}}\,\bm{\xi}^{r}_{i}.$ (1b)
In the above equations, $\mu$ is mobility, $D_{t}$ and $D_{r}$, respectively, are translational and rotational diffusion constants of the particle, while $\bm{\xi}^{t}$ and $\bm{\xi}^{r}$ are white noises with zero mean and unit variances. The constant $\chi_{r}$ is taken to be positive so that the particles rotate away from each other (chemo-repulsive). The force on the $i$th particle is given as: $\mathbf{F}_{i}^{b}=-{\partial U}/{\partial\mathbf{r}_{i}}$, while we have: $U=\sum^{N-1}_{i=1}\mathcal{U}^{b}(\mathbf{r}_{i},\mathbf{r}_{i+1})$. Here, $\mathcal{U}^{b}(\mathbf{r}_{i},\mathbf{r}_{j})=k_{b}\left(r_{ij}-r_{0}\right)^{2}$ is a harmonic potential of stiffness $k_{b}$ and natural length $r_{0}$ which holds the chain together, while $r_{ij}=|\mathbf{r}_{i}-\mathbf{r}_{j}|$.
@@
$\theta_{i}(0)$ $=0,\qquad x_{i}(0)=0,\qquad\forall i,\qquad\qquad$ (3a)
$y_{i+1}(0)$ $=y_{i}(0)+2b,\qquad\qquad\,\,\,1\leq i\leq N-1$ (3b)
where $b$ is the radius of each colloid. Note that $r_{0}=2b$. All numerical results in this paper are generated with an explicit Euler-Maruyama integrator with time step $dt=0.01$ and simulation time $T=7.2\times 10^{5}$. For all simulations here (unless stated otherwise) $b=1$, $r_{0}=2b$, $c_{0}=1$, $D_{c}=1$, $\hat{\mu}=\mu k_{b}=91.6$, and $v_{s}=1$. $D_{r}=0$ for all sections, with the exception of Sec. V.3 where values are explicitly stated. Wherever realization average is performed, we use $50$ realizations, with the exception of $10$ used in Appendix .1.
@@ III Dynamical regimes
The dimensionless number $\Lambda$ quantifies the competition between deterministic rotations (due to phoretic interactions) and deterministic propulsion. The dimensionless number $\tilde{\Lambda}$ quantifies deterministic and random motion in the positional sector. A variant on the above model, proposed in section II, was introduced in [25] and studied further in [29]. In both these papers, we had considered the role of trail created by the particles. Here, we first assume instantaneous chemical interactions, and discuss the effect of trail-mediated history later (in Appendix .1). Additionally, we note that the previous works ignored the role of translational noise $D_{t}$. In [29], it was found that a competition between $\tau_{r}$ and the chemical diffusivity of trails leads to the formation of a stable C-shape chain.
@@
It is informative to study the model described above in terms of the two dimensionless numbers in Eq. (5). In Fig. 1, we present a phase diagram in the plane of $\Lambda$ and $\tilde{\Lambda}$. We find that the C-shape is sustained for a selected range of dimensionless numbers $\Lambda$ and $\tilde{\Lambda}$ [25, 29]. Let us construct in addition a fluctuating length scale $l_{f}=\frac{D_{t}}{v_{s}}$ and a correlation length scale $l_{C}=\frac{\chi_{r}}{bv_{s}}$ to aid with the analysis. First, in the limit of $\Lambda\gtrsim 10^{2},\tilde{\Lambda}\gtrsim 10^{1}$, there exists insufficient rotations and the chain remains flat within the simulation time scale [29]. This corresponds to the region of extremely short-wavelength fluctuations ($l_{f}<b$) and small deterministic correlation lengths ($l_{C}<10^{-3}b$). This straight-chain phase is labelled as Phase I (see Fig. 1(b) top left). Next, for regions $\tilde{\Lambda}\lesssim 10^{0}$, we find a disordered phase with effectively zero positional order in the chain. Here we find that $l_{f}>b$, thus we conclude that long-wavelength fluctuations of greater than the monomer size are not supported by our colloidal chain. This disordered phase is labelled as Phase II (Fig. 1(b) top right). Further, focusing in the regime $\tilde{\Lambda}>10^{1}$ and $\Lambda<10^{-3}$ corresponds to the case where deterministic rotations are instantaneous across the system ($\tau_{f}<<\tau$), such that the chemical interactions are instantaneously correlated. Here one has that $l_{C}\gtrsim 10^{3}b$, thus correlations much larger than the system size. The dynamics of the chain displays “frustrated” behavior, not attaining any non-trivial spatial structure as there is not enough time to respond to instantaneous chemical gradients. This frustrated phase is labelled as Phase III (Fig. 1(c) bottom left). The remaining region corresponds to that of the C-shape (Phase IV). We thus conclude that the C-shape corresponds to regimes where the correlation lengths $l_{C}$ are of the order of the chain length, which is in addition stable to short-wavelength fluctuations (of typical distance less that one monomer radius). It is this dynamical steady-state (DSS) phase that we study in the rest of this paper.
@@
We also note that, within the C-shape phase, its existence as a DSS depends on the choice of initial conditions of Eq. (3). Deviating from these does not render a universality of the DSS, unless one includes the existence of chemical trails in the model [25, 29]. This scenario is further discussed in Appendix .1. If one deviates from the purely chemo-repulsive scenario considered here, the presence of trails is also necessary for other non-trivial DSSs, such as a swimming/undulating state. However, various structural SSs (e.g. designer crystallites that break handedness symmetry) are possible via the instantaneous chemical deposition model considered here, though one would have to significantly alter from the fully chemo-repulsive scenario, and incorporate mixed (anti-symmetric and/or non-reciprocal) interactions [29].
@@ III.1 Characteristics of the C-shape
    Finite mean curvature of the chain
@@
As explained in Appendix .7, there are various other DSSs that meet some of the above requirements but not all. For instance the aforementioned Phase I DSS satisfies only (i), (iii) and (iv). The C-shape is unique in that it is the only DSS that meets requirements (i)-(iv), with its propulsion perpendicular to the direction of the body axis. It is in this specific phase in which the fluctuations are computed.
@@
Table 1: Table of exponents defined in equations (6-11) of section IV. Note that $z$ is defined such that $\beta^{(2)}$ is in the denominator for the first and third rows, whereas $\beta^{(1)}$ is in the denominator for the others (orientational and locally flat exponents). For the last two columns, $z$ is taken to be $z^{\prime}$ implicitly.
@@ III.2 C-shape selection
In Phase IV, as mentioned previously, the C-shape is always attained under the initial conditions of 3. We note that in such an autophoretic chain with chemical trails this DSS is the universal solution regardless of initial conditions [25, 29] (see Appendix .1)). However, it is instructive to ask to what extent the C-shape is attained under different initial conditions.
To this end, we apply a hysteretic protocol by varying $\Lambda$ and carrying on the final state of a given simulation to the initial conditions of the simulation at the next time step. The order parameter used to identify this DSS is the absolute value of the curvature, $|\kappa|$ defined in Eq. (A15). Two separate cases are presented in Fig. 1(c), corresponding to decrease of $\Lambda$ starting from Phase I (the stiff chain; Fig. 1(c), top), and increase of $\Lambda$ starting from Phase IV (the C-shape; Fig. 1(c), bottom). We observe from the former that the stiff-to-C transition is irreversible. Thus, a C-shape DSS cannot lose its spatial curvature via a decrease in self-phoretic torques. From the latter we see that the C-to-stiff transition is hence not permitted; both forward and backward sweeps of the protocol do not change the overall curvature. We also note that we have separately varied $\tilde{\Lambda}$ to and from Phase IV (not shown here); the results of these indicate that selected random configurations of Phase II may indeed transition to IV, but again there is no clear reliable reversibility of transition to define a hysteresis loop. This irreversibility of the first transition implies that such a hysteresis curve and hence a mapping to known classes of equilibrium or dynamical phase transitions is not well defined (indeed one may have to devise an alternate order parameter); we leave the exact characterization of the nature of the transition to and from the C-shape for future work.
@@ IV Interface fluctuations
For the C-shape chain propelling in the positive $x$-direction, one can define a height function $h(y)$ along the chain. In particular $h$ corresponds to the $x$-displacement of each monomer from the mean height. An example of its temporal evolution is shown in Fig. 1(d). One can thus choose a particular length segment along the chain and compute the height fluctuations, and measure its root-mean-squared value
@@
The quantities of interest are both the dynamical ($W_{h}(t\ll\infty)$ and $W_{\theta}(t\ll\infty)$) and steady-state ($W^{\mathrm{ss}}_{h}(t\rightarrow\infty)$ and $W^{\mathrm{ss}}_{\theta}(t\rightarrow\infty)$) properties of these fluctuations. These are:
$W_{h}(t)$ $\sim t^{\beta_{h}},\qquad W^{\mathrm{ss}}_{h}\sim N^{\alpha_{h}},$ (8a)
$W_{\theta}(t)$ $\sim t^{\beta_{\theta}},\qquad W^{\mathrm{ss}}_{\theta}\sim N^{\alpha_{\theta}}.$ (8b)
@@
To compute the quantities $W_{h}(t)$ and $W_{\theta}(t)$ [Eq. (6) and (7)] respectively, we discuss two possibilities. There are two separate segments about which fluctuations are measurable. Let us label the length of this segment to be $N^{\prime}$. The first obvious choice of the segment is $N^{\prime}=N$, corresponding to the average over the entire C-shape. The second corresponds to that of a locally flat region; here $N^{\prime}<<N$, which is taken to be sufficiently far from the edges of the chain. Though established results exist for fluctuations about curved segments of an interface [9], these typically correspond to that of a (subset of a) circular interface, thus different from that of the C-shape (see Appendix .5). For the latter, it is to be noted that fluctuations of growing (circular) interfaces are typically measured in this regime [36, 37] ; the scaling properties in this region are almost always equal to that of globally measured quantities - for instance in [16] for a flat surface and [36] for a curved surface. We discuss the specific instances of equivalences and/or differences for the C-shape below.
All results presented in this paper are computed for values of $N\in[32,64,100,128,192,200,256,312]$, with curves for specific $N$ displayed (and labeled) in each respective figure. The results for the incorporation of rotational noise are computed for $N\in[64,128,192,256,312]$. The choices for $N^{\prime}$ are explicitly stated in the respective sections.
@@ V Results
A summary of the results of this paper is presented in Table 1, wherein the respective Figures are referenced. It is to be noted that the value of $\beta$ for each respective figure is obtained through the best possible collapse of the scaling relations 9 and 11 (by eye). The values for $\alpha$ on the other hand are obtained by a least-squares fit on the steady state values of $W$ for each system size. For each plot of $W$, the fitting windows used for the in-figure $\beta$ line of best fit is mentioned in the caption.
@@ V.1 Height fluctuations
From (14), (15) and Fig. 2, we find that the regimes (iii) and (iv) exhibit a clear FV scaling with a roughness exponent $\alpha_{h}$ and a dynamic exponent $z_{h}=\frac{\alpha_{h}}{\beta_{h}}\approx 0.52$. To the best of our knowledge, this set of exponents does not correspond to any previously reported universality class. We thus interpret this as a novel “C-shape roughness” scaling, distinct to those found for either circular or flat interfaces [36, 15]. The super-ballistic exponent can be appreciated by noting that as the C-shape formation takes place, displacements of each monomer from the mean is enhanced in either direction (along the x-axis) along the chain. These squared deviations from the mean height are in addition symmetric along the y-axis. $l_{c}$ is clearly enhanced due to this symmetry. It is to be noted that in a conventional ballistic scaling one simply has uni-directional deviations along the propulsion axis; in our system this is applicable to the mean $h(y)$; clearly $l_{c}$ in this case is smaller than that for a symmetric C-shape. We suggest that these combined effects give rise to the super-ballistic growth exponent $\beta_{h}$. The emergence of a super-ballistic exponent can be rationalized via a linearised analysis of the EOM 1. This is presented in Appendix .3. It is to be noted that varying $\tau_{r}$ changes the time at which super-ballistic growth sets in (exponent and roughness is unchanged). The roughness exponent $\alpha_{h}$, on the other hand, indicates the relative growth of height fluctuations when $l_{c}$ spans the system size. Here, the system is at $\sim 10^{2}\tau_{f}$ and the C-shape is fully formed. We note that the value we report for $\alpha_{h}\approx 0.9$ is greater than the standard roughness exponent encountered in dynamics without height conservation ($\alpha_{h}\approx 0.5$[34, 35] and smaller than that in models with height conservation ($\alpha_{h}\approx 1.5$)[17]. We elaborate on these findings further in Sec. VI.